OpenStudy (anonymous):

In a problem it says that x is "at most –5." What does this mean?

OpenStudy (nikato):

that means x is less than or equal to -5. it can not be greater than -5

OpenStudy (nikato):

\[x \le-5\]

OpenStudy (anonymous):

X<=-5 like t

OpenStudy (anonymous):

this

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

yea, thats right

OpenStudy (anonymous):

An engine piston has a diameter of 9.2 cm, with a tolerance of 0.01 cm. Which absolute value inequality represents the range of values for the diameter of the piston?

OpenStudy (anonymous):

can you help me in this one @nikato

OpenStudy (nikato):

so with the same thing, i think. the piston thing has to be within 0.01 difference from 9.2

OpenStudy (nikato):

so which 2 numbers are 0.1 greater than 9,2 and 0.1 less than 9.2

OpenStudy (anonymous):

9,2

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

|x – 0.01| <= 9.2

OpenStudy (nikato):

i think its the other way around. switch the 9.2 and 0.01

OpenStudy (anonymous):

|x – 9.2| <= 00.1

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

correct